OpenStudy (anonymous):

PLEASE HELP I HAVE A HUGE TEST TOMORROW a baseball player puts a baseball on a batting tee. the baseball bat is able to deliver 4.0 * 10^5 Watts of power and is in contact with the ball for 0.70 milliseconds and a distance of 1.4 centimeters. the mass of the ball is 145 grams. what is the net force applied to the ball and what is its average acceleration? net force is supposed to be 2.0 * 10^4 N and acceleration is supposed to be 1.4 * 10^5 m/s^2 but i cant figure out how

OpenStudy (anonymous):

...I cannot help you with your question. BUT I can help you obtain a valuable thing... all you ave to do, is reply to your own question ,and I may help you recieve a medal.

OpenStudy (anonymous):

ur wasting your time kid

OpenStudy (anonymous):

how so? I believe you just received a medal no?

OpenStudy (anonymous):

i dont need a medal... i need answers to my question... im here to study not fool around

OpenStudy (anonymous):

eh, it's better than me coming in and saying hi, and leaving. at least you get something, although it's not rly useful.

OpenStudy (anonymous):

ur just wasting my time too

OpenStudy (anonymous):

your not getting an answer any faster if i'm not here.

OpenStudy (anonymous):

im trying other sources

OpenStudy (anonymous):

\(\Sigma P =\frac{Fs}{t}\) where P=4.0 * 10^5 , s=0.14 m, t=7*10^-4 s plug it all in. do you need further help?

OpenStudy (anonymous):

thanks man... but my dad already helped me on this... :D

OpenStudy (anonymous):

lol You have a smart dad :) you're welcome